# Generative AI for Cybersecurity of Energy Management Systems: Methods, Challenges, and Future Directions

### v1

@@ Abstract
This paper elaborates on an extensive security framework specifically designed for energy management systems (EMSs), which effectively tackles the dynamic environment of cybersecurity vulnerabilities and/or system problems (SPs), accomplished through the incorporation of novel methodologies. A comprehensive multi-point attack/error model is initially proposed to systematically identify vulnerabilities throughout the entire EMS data processing pipeline, including post state estimation (SE) stealth attacks, EMS database manipulation, and human-machine interface (HMI) display corruption according to the real-time database (RTDB) storage. This framework acknowledges the interconnected nature of modern attack vectors, which utilize various phases of supervisory control and data acquisition (SCADA) data flow. Then, generative AI (GenAI)-based anomaly detection systems (ADSs) for EMSs are proposed for the first time in the power system domain to handle the scenarios. Further, a set-of-mark generative intelligence (SoM-GI) framework, which leverages multimodal analysis by integrating visual markers with rules considering the GenAI capabilities, is suggested to overcome inherent spatial reasoning limitations. The SoM-GI methodology employs systematic visual indicators to enable accurate interpretation of segmented HMI displays and detect visual anomalies that numerical methods fail to identify. Validation on the IEEE 14-Bus system shows the framework’s effectiveness across scenarios, while visual analysis identifies inconsistencies. This integrated approach combines numerical analysis with visual pattern recognition and linguistic rules to protect against cyber threats and system errors.
Keywords Anomaly detection  $\cdot$ Attack  $\cdot$ Energy management systems  $\cdot$ Generative AI  $\cdot$ Multi-modal Analysis  $\cdot$ Set-of-Mark Generative Intelligence.
1 Introduction
@@
The accelerated digitization of power systems has revolutionized EMSs into complex cyber-physical architectures capable of coordinating electricity delivery across expansive grids. Modern EMS frameworks integrate SCADA systems with cutting-edge computational algorithms, promoting a complex network in which field equipment such as remote terminal units (RTUs) and phasor measurement units (PMUs) transmit synchronized, real-time data that enables enhanced monitoring and control of grid performance [1, 2, 3]. While this integration brings substantial operational efficiencies and enhanced operational visibility, it also expands the system’s vulnerability footprint, exposing multiple attack vectors within the data acquisition and processing sequence. The convergence of operational technology (OT) and information technology (IT) domains necessitates a reconsideration of traditional cybersecurity approaches. SE is crucial in power systems, offering operators a real-time overview of the system’s condition to maintain reliability and control [4]. However, the coupling of power systems with cyber infrastructure increases their exposure to complex cyber-attacks. Among these, false data injection (FDI) attacks present a particularly serious threat by manipulating measurement data to lead system operators toward incorrect control actions while evading traditional bad data detection (BDD) methods [5]. Also, stealth attacks represent a sophisticated subset of FDI attacks. These attacks are mathematically designed to bypass traditional BDD mechanisms by ensuring that the measurement residuals remain practically unchanged from normal conditions [6]. Hence, BDD techniques such as $\chi^{2}$-test primarily based on SE residuals are increasingly insufficient, particularly against sophisticated stealth attacks and FDI strategies. These stealthy attacks can manipulate PMU and SCADA data while maintaining consistency with system models, evading conventional BDD filters [7, 8]. In parallel, attackers might manipulate data within HMIs to introduce misleading display changes that mislead operators, all the while not necessitating the modification of actual physical measurements. Given the critical infrastructure, modern power grids necessitate robust mechanisms capable of identifying both mathematically stealthy manipulations and operator-level display tampering. Recent literature emphasizes hybrid detection approaches that synthesize physical-model awareness, data-driven analytics, and temporal behavioral models to detect anomalies across multi-layered attack vectors [9]. In accordance with this, testbed-based experiments and co-simulation studies emphasize the importance of AI and multi-modal detection approaches to limit false positives (FPs) while improving sensitivity to complex attacks [10, 11, 12].
1.1 Problem Statement
@@
The evolving cybersecurity threats targeting EMSs represent a new class of challenges that exceed the capabilities of traditional defenses originally designed for isolated OT domains [10]. As EMS architectures increasingly integrate both cyber and physical layers, malicious actors are now capable of executing coordinated attacks that span the entire data processing pipeline from the collection of measurements through the SE stage, and finally to the HMI screen. An alarming feature of such breaches is their ability to maintain statistical validity, enabling them to evade conventional BDD techniques such as those relying on $\chi^{2}$ hypothesis testing. These attack pathways are diverse and sophisticated, such that some are stealth attacks or may involve tampering with SE outputs after validation but before storage in historical databases, effectively corrupting trusted data storage systems. Others may directly target the EMS databases, compromising both operational and historical data integrity. Also, some attacks on HMIs can happen, which can subtly manipulate visual representations such as circuit breaker (CB) statuses or voltage levels without modifying the core numerical calculations, thereby misleading operators and delaying response times despite accurate SE outputs [10]. The complexity of power systems, intensified by the integration of distributed energy resources (DERs), further increases these vulnerabilities. Within such dynamic conditions, malicious data injections can be designed to imitate normal operations, potentially leading the system into inefficient or unstable operating states without timely identification.
@@
Also, traditional anomaly detection (AD) mechanisms such as the SE process focus on numerical inconsistencies and threshold-based alarms, which are often insufficient when attackers manipulate the visual elements of SCADA displays. For instance, falsified visual indicators may not breach statistical thresholds but can still mislead human operators, especially during high-stress operational conditions. Although recent advances in GenAI offer promising opportunities for pattern recognition and AD processes, they present distinctive challenges. Many generative models still lack robust spatial reasoning capabilities and often require explicit guidance or structured prompts to accurately interpret the content of segmented or context-rich HMI displays. The absence of integrated detection frameworks which simultaneously assess numerical validity, visual consistency, and semantic rule coherence creates substantial gaps in current EMS cybersecurity strategies. These gaps provide opportunities for sophisticated attackers to exploit the interfaces between detection layers and human perception, particularly in scenarios requiring rapid and confident operator decision-making. Addressing these vulnerabilities requires a framework shift toward multi-modal, intelligent security solutions capable of bridging these aspects of grid operations [13].
1.2 Research Objectives
@@
1.3 Related Work
Generative pre-trained transformers (GPTs) can enhance the system diagnosis (SD) accuracy beyond traditional machine learning (ML) models and BDDs through contextual processing and adaptability. Unlike ML models requiring extensive pattern training, GPTs can simultaneously analyze data, historical patterns, and operator inputs. Their NLP capabilities enable the integration of operator logs with numerical data, leading to more accurate SD processes in EMSs. They can also adapt to new system configurations without complete retraining, making them more efficient for evolving power network topologies while maintaining high diagnostic accuracy during critical situations [14, 15, 16, 17]. Furthermore, it is challenging for BDDs to detect stealth attacks as they evade detectors as well as the unexpected scenarios. Hence, a summary of some related research is given in Table 1.
Table 1: A literature survey on the AD process in EMSs.
@@ 1.3 Related Work
Ashrafuzzaman et al. [20] introduced a data-driven ensemble ML framework aimed at the detection of stealthy FDI attacks within smart grids by employing classification algorithms. Although this approach enhances the AD process by mitigating the complexities associated with high-dimensional data, it faces significant challenges, particularly the requirement for extensive labeled datasets and the frequency of high FP rates in unsupervised scenarios. An approach addressing stealth sensor and actuator attacks under resource constraints on discrete event systems using supervisory control was developed by He et al. [24]. They introduced combined vulnerability and established efficient algorithmic techniques. However, their approach is constrained by the computational complexity of modeling extensive discrete event system scenarios. Guo et al. [21] introduced an event-driven stealthy FDI attack strategy against remote SE systems. Their method dynamically initiates attacks based on real-time residuals to deteriorate system performance optimally. Despite its effectiveness in resource-limited contexts, it depends on accurate real-time residual computations, which could be challenging in noisy environments.
@@
A perturbation strategy for defending against FDI attacks in IoT-based smart grids was presented by Zhang et al. [22]. They demonstrated theoretically and numerically that inappropriate selection of branches for reactance perturbation could compromise defense effectiveness. However, their enhanced strategy requires detailed prior topology analysis, potentially limiting its scalability to larger systems. Guo et al. [23] introduced a residual-based stealthy FDI attack for multi-sensor estimation systems, highlighting the critical selection of sensors under resource constraints to optimize degradation. Their approach leverages historical and current residuals to enhance attack impact but demands significant computational resources to solve optimization problems at each step. Zhou et al. [25] investigated optimal FDI attacks against partially secured remote SE systems by formulating optimization problems to maximize SE errors. They developed robust detection and resilient estimation strategies; yet the practicality of their methods could be affected by the complexity of real-world implementation and the necessity of secured communication channels. A framework utilizing ML algorithms to approximate the PF analysis was developed by Falconer et al. [18]. This approach supported the resolution of complex PF problems that incorporated unit commitment and security restrictions. However, their model faces challenges stemming from scalability issues associated with fully connected networks as the system size increases. Additionally, CNN models demonstrate restricted predictive accuracy due to their dependence on convolutions if there are anomalies. Hu et al. [26] introduced the state deviation index for the diagnosis of FDI attacks and sudden load changes, demonstrating its effectiveness in IEEE 14-bus and 30-bus systems. However, the algorithm exhibited limited adaptability when faced with different errors. This suggests the need for an approach in which indicators are carefully designed to reflect unique properties of specific anomalies. The absence of such tailoring risks compromising the accuracy of evaluation metrics, presenting significant FP and false negative (FN) results. A multi-variable long short-term memory autoencoder (LSTM-AE) had been formulated for an SD process by Sarker et al. [27]. Their model successfully diagnosed SPs associated with errors, validated through a 123-bus unbalanced distribution network. However, according to sophisticated mathematical modeling of the PF analysis considering the SD process, developing the model with different abnormal scenarios in ML algorithms is challenging. Since they jeopardize the accuracy of the proposed algorithm, which can be time-consuming, they need more effort. Furthermore, a graphical user interface (GUI) of an EMS could show abnormal information due to system errors or bugs. These errors are not easy for SCADA control room engineers to diagnose due to the huge volume of information. These issues cannot be managed by the SE process and need adaptive solutions. A novel FDI diagnostic method using LSTM-AE and CNN-AE with an unsupervised learning approach was introduced in [28], avoiding the need for anomalous data during training. It also proposed an LSTM variational AE-based reconstruction method to maintain stability by closely replicating the original data from anomalous data. Despite this, the reliance on unsupervised learning presents challenges in certain scenarios. Hence, zero-day attacks need retraining of ML algorithms, which is time-intensive. Also, collecting all unknown errors can take much effort [29, 30]. Mukherjee [19] introduced an approach to identify FDI attacks using a multi-label classification framework. This method leveraged conventional bad-data detectors to enhance measurement accuracy and diagnose unstructured attacks. Accordingly, this model-free strategy required no prior knowledge of grid or attack vectors, making it a highly effective solution for the FDI diagnosis. Nevertheless, a consideration of an FDI attack cannot merely show the comprehensiveness of this algorithm. Also, there are some malfunctions that cannot be properly diagnosed by ML models. A retraining on new attacks and the diagnosis of other malfunctions (e.g., a CB can be opened during a fault in the normal operation; however, this could be because of inaccurate communication that sends the status of the CB to the SCADA room) are challenging for ML techniques [31, 32, 33].
1.4 Contributions
@@
  • Multi-Point Attack Detection Framework along with GenAI-based AD Processes: A GenAI-based ADS model developed that revolutionizes how vulnerabilities are identified and addressed within the EMS data processing pipeline based on the multi-point attacks/errors. In contrast to conventional approaches that focus on isolated detection points, this framework recognizes that sophisticated cyber-attacks exploit multiple stages of the SCADA data flow. The model specifically targets three critical vulnerability points including stealth attacks which can evade the BDDs, EMS database manipulation incidents (particularly FDI attacks following the SE process), and the HMI display corruption by manipulating the RTDB. Through implementation on the IEEE 14-bus system, the framework’s capability demonstrates to detect attack patterns including state vector manipulations, topology corruption schemes, and display RE attacks, all while preserving mathematical consistency to bypass conventional BDD algorithms (e.g., $\chi^{2}$ method). This integrated approach represents a transformative approach in the EMS security, acknowledging that modern cyber threats require comprehensive monitoring across the entire data processing flow rather than relying on merely the mathematical consistency.
@@
  • Generative Intelligence-Enhanced Multimodal Analysis Framework: Addressing the integration challenges and anomaly identification requirements, an innovative SoM-GI framework is proposed. This contribution bridges the gap between traditional numerical SE techniques and optimal PF results and emerging visual AD capabilities. The framework leverages the power of GenAI, specifically enhanced with engineered visual markers and indicators in additions to the rules, to enable simultaneous analysis of both image and textual data within the EMS environment. The SoM-GI methodology overcomes inherent spatial reasoning limitations in current GenAI systems by implementing systematic visual indicators, including CB status markers, directional transmission line indicators, and CP identifiers. These markers guide the AI’s interpretation of segmented HMI displays, enabling accurate detection of visual anomalies inconsistencies and falsifications that often bypass numerical detection methods. Validation results confirm the framework’s ability to identify sophisticated attacks that manipulate display segments while maintaining consistency in core data structures. This novel approach establishes a new paradigm for comprehensive security monitoring in power systems, integrating visual pattern recognition with linguistic rule processing. By standardizing the fusion of multimodal analysis techniques, the framework ensures that visual-based ADSs can be incorporated into existing EMS architectures without disrupting critical operational processes, thereby providing a practical solution.
1.5 Paper Structure

### v2

@@ Abstract
This paper elaborates on an extensive security framework specifically designed for energy management systems (EMSs), which effectively tackles the dynamic environment of cybersecurity vulnerabilities and/or system problems (SPs), accomplished through the incorporation of novel methodologies. A comprehensive multi-point attack/error model is initially proposed to systematically identify vulnerabilities throughout the entire EMS data processing pipeline, including post state estimation (SE) stealth attacks, EMS database manipulation, and human-machine interface (HMI) display corruption according to the real-time database (RTDB) storage. This framework acknowledges the interconnected nature of modern attack vectors, which utilize various phases of supervisory control and data acquisition (SCADA) data flow. Then, generative artificial intelligence (GenAI)-based anomaly detection systems (ADSs) for EMSs are proposed for the first time in the power system domain to handle the scenarios. Further, a set-of-mark generative intelligence (SoM-GI) framework, which leverages multimodal analysis by integrating visual markers with rules considering the GenAI capabilities, is suggested to overcome inherent spatial reasoning limitations. The SoM-GI methodology employs systematic visual indicators to enable accurate interpretation of segmented HMI displays and detect visual anomalies that numerical methods fail to identify. Validation on the IEEE 14-Bus system shows the framework’s effectiveness across scenarios, while visual analysis identifies inconsistencies. This integrated approach combines numerical analysis with visual pattern recognition and linguistic rules to protect against cyber threats and system errors.
Index Terms:
Anomaly Detection, Attack, EMS, GenAI, Multi-modal Analysis, SoM-GI.
[…]
=-15pt
I Introduction
@@
The accelerated digitization of power systems has revolutionized EMSs into complex cyber-physical architectures capable of coordinating electricity delivery across expansive grids. Modern EMS frameworks integrate SCADA systems with cutting-edge computational algorithms, promoting a complex network in which field equipment such as remote terminal units (RTUs) and phasor measurement units (PMUs) transmit synchronized, real-time data that enables enhanced monitoring and control of grid performance [1, 2, 3]. While this integration brings substantial operational efficiencies and enhanced operational visibility, it also expands the system’s vulnerability footprint, exposing multiple attack vectors within the data acquisition and processing sequence. The convergence of operational technology (OT) and information technology (IT) domains necessitates a reconsideration of traditional cybersecurity approaches. SE is crucial in power systems, offering operators a real-time overview of the system’s condition to maintain reliability and control [4]. However, the coupling of power systems with cyber infrastructure increases their exposure to complex cyber-attacks. Among these, false data injection (FDI) attacks present a particularly serious threat by manipulating measurement data to lead system operators toward incorrect control actions while evading traditional bad data detection (BDD) methods [5]. Also, stealth attacks represent a sophisticated subset of FDI attacks. These attacks are mathematically designed to bypass traditional BDD mechanisms by ensuring that the measurement residuals remain practically unchanged from normal conditions [6]. Hence, BDD techniques such as $\chi^{2}$-test, primarily based on SE residuals, are increasingly insufficient, particularly against sophisticated stealth attacks and FDI strategies. These stealthy attacks can manipulate PMU and SCADA data while maintaining consistency with system models, evading conventional BDD filters [7, 8]. In parallel, attackers might manipulate data within HMIs to introduce misleading display changes that mislead operators, all the while not necessitating the modification of actual physical measurements. Given the critical infrastructure, modern power grids necessitate robust mechanisms capable of identifying both mathematically stealthy manipulations and operator-level display tampering. Recent literature emphasizes hybrid detection approaches that synthesize physical-model awareness, data-driven analytics, and temporal behavioral models to detect anomalies across multi-layered attack vectors [9]. In accordance with this, testbed-based experiments and co-simulation studies emphasize the importance of AI and multi-modal detection approaches to limit false positives (FPs) while improving sensitivity to complex attacks [10, 11, 12].
I-A Problem Statement
@@
The evolving cybersecurity threats targeting EMSs represent a new class of challenges that exceed the capabilities of traditional defenses originally designed for isolated OT domains [10]. As EMS architectures increasingly integrate both cyber and physical layers, malicious actors are now capable of executing coordinated attacks that span the entire data processing pipeline, from the collection of measurements through the SE stage and finally to the HMI screen. An alarming feature of such breaches is their ability to maintain statistical validity, enabling them to evade conventional BDD techniques such as those relying on $\chi^{2}$ hypothesis testing. These attack pathways are diverse and sophisticated, such that some are stealth attacks or may involve tampering with SE outputs after validation but before storage in historical databases, effectively corrupting trusted data storage systems. Others may directly target the EMS databases, compromising both operational and historical data integrity. Also, some attacks on HMIs can occur, which can subtly manipulate visual representations such as circuit breaker (CB) statuses or voltage levels without modifying the core numerical calculations, thereby misleading operators and delaying response times despite accurate SE outputs [10]. The complexity of power systems, intensified by the integration of distributed energy resources (DERs), further increases these vulnerabilities. Within such dynamic conditions, malicious data injections can be designed to imitate normal operations, potentially leading the system into inefficient or unstable operating states without timely identification.
@@
Also, traditional anomaly detection (AD) mechanisms, such as the SE process, focus on numerical inconsistencies and threshold-based alarms, which are often insufficient when attackers manipulate the visual elements of SCADA displays. For instance, falsified visual indicators may not breach statistical thresholds but can still mislead human operators, especially during high-stress operational conditions. Although recent advances in GenAI offer promising opportunities for pattern recognition and AD processes, they present distinctive challenges. Many generative models still lack robust spatial reasoning capabilities and often require explicit guidance or structured prompts to accurately interpret the content of segmented or context-rich HMI displays. The absence of integrated detection frameworks that simultaneously assess numerical validity, visual consistency, and semantic rule coherence creates substantial gaps in current EMS cybersecurity strategies. These gaps provide opportunities for sophisticated attackers to exploit the interfaces between detection layers and human perception, particularly in scenarios requiring rapid and confident operator decision-making. Addressing these vulnerabilities requires a framework shift toward multi-modal, intelligent security solutions capable of bridging these aspects of grid operations [13].
I-B Research Objectives
@@
I-C Related Work
Generative pre-trained transformers (GPTs) can enhance the system diagnosis (SD) accuracy beyond traditional machine learning (ML) and deep learning (DL) models and BDDs through contextual processing and adaptability. Unlike ML models requiring extensive pattern training, GPTs can simultaneously analyze data, historical patterns, and operator inputs. Their NLP capabilities enable the integration of operator logs with numerical data, leading to more accurate SD processes in EMSs. They can also adapt to new system configurations without complete retraining, making them more efficient for evolving power network topologies while maintaining high diagnostic accuracy during critical situations [14, 15, 16, 17]. Furthermore, it is challenging for BDDs to detect stealth attacks as they evade detectors as well as unexpected scenarios. Thus, Table I provides a concise review of relevant studies.
TABLE I: A literature survey on the AD process in EMSs.
@@ I-C Related Work
Ashrafuzzaman et al. [20] introduced a data-driven ensemble ML framework aimed at the detection of stealthy FDI attacks within smart grids by employing classification algorithms. Although this approach enhances the AD process by mitigating the complexities associated with high-dimensional data, it faces significant challenges, particularly the requirement for extensive labeled datasets and the frequency of high FP rates in unsupervised scenarios. An approach addressing stealth sensor and actuator attacks under resource constraints on discrete event systems using supervisory control was developed by He et al. [24]. They introduced combined and efficient vulnerability and algorithmic techniques. However, their approach is constrained by the computational complexity of modeling extensive discrete event system scenarios. Guo et al. [21] introduced an event-driven stealthy FDI attack strategy against remote SE systems. Their method dynamically initiates attacks based on real-time residuals to deteriorate system performance optimally. Despite its effectiveness in resource-limited contexts, it depends on accurate real-time residual computations, which could be challenging in noisy environments.
@@
Lee et al. [25] proposed a unified industrial large knowledge model framework, emphasizing how domain-specific knowledge can be integrated with large language models (LLMs) to address complex industrial challenges. Their framework introduces the concept of combining human-interpretable data with structured machine-generated data through LLM-based semantic understanding. The authors demonstrated that LLMs can serve as domain experts by processing diverse industrial data and generating actionable insights, though they noted challenges in reliability and privacy that require specialized training approaches. Building on industrial LLM applications, Aberbach et al. [26] investigated how LLMs can enhance smart grid resilience and efficiency by processing unstructured text-based information alongside traditional grid data. Their work highlighted that LLMs excel at handling diverse data types including operator queries, maintenance logs, and regulatory documents. They emphasized the importance of combining LLMs with specialized algorithms for grid-specific tasks, suggesting a hybrid approach similar to the integration of prompt engineering with meta-learning techniques, without multimodal analysis considering the analysis of HMI screen and visual interpretation of power system components. Further, Li et al. [27] introduced a graph transformer-based vision-language model for industrial AD processes. Their research demonstrated that aligning visual features with textual descriptions through prompt engineering significantly improves AD accuracy across different product categories. The study showed that multi-level domain adaptation enables models to capture semantic consistency of anomalous features at various scales, achieving strong performance even with limited training data. However, their approach focused primarily on visual anomalies in manufacturing rather than cyber-physical attacks in smart grids. More directly related to grid security, Shen et al. [28] developed an LLM-based security situation awareness framework specifically for smart grids. Their approach transformed diverse grid data into structured text prompts and utilized LLM reasoning capabilities to detect threats and forecast security states. The research demonstrated that LLMs could process electrical parameters, meteorological data, and social factors in a unified semantic space, achieving improved prediction accuracy through multimodal fusion. Nevertheless, their work concentrated on general security awareness rather than addressing the specific challenge of zero-day attack detection in federated environments. Addressing privacy concerns in distributed grid systems, Dasgupta and Mitra [29] proposed a federated zero-shot learning framework that leverages LLMs to generate client-specific semantic embeddings for intrusion detection. Their approach demonstrated how LLMs could create varied textual descriptions of attacks for each client, reducing the risk of inference attacks while maintaining detection accuracy. By combining federated learning with zero-shot capabilities, they showed that models could detect unknown attacks without sharing sensitive grid data. However, their methodology differs from the proposed approach in that they focused specifically on intrusion detection for network-based attacks rather than the broader operational AD and forecasting tasks considering the visual interpretation and power flow modeling. A perturbation strategy for defending against FDI attacks in IoT-based smart grids was presented by Zhang et al. [22]. They demonstrated theoretically and numerically that inappropriate selection of branches for reactance perturbation could compromise defense effectiveness. However, their enhanced strategy requires detailed prior topology analysis, potentially limiting its scalability to larger systems. Guo et al. [23] introduced a residual-based stealthy FDI attack for multi-sensor estimation systems, highlighting the critical selection of sensors under resource constraints to optimize degradation. Their approach leverages historical and current residuals to enhance attack impact but demands significant computational resources to solve optimization problems at each step. Zhou et al. [30] investigated optimal FDI attacks against partially secured remote SE systems by formulating optimization problems to maximize SE errors. They developed strong detection and estimation strategies, but real-world complexity and the need for secure communication may impact practicality. A framework utilizing ML algorithms to approximate the PF analysis was developed by Falconer et al. [18]. This approach supported the resolution of complex PF problems that incorporated unit commitment and security restrictions. However, their model faces challenges stemming from scalability issues associated with fully connected networks as the system size increases. Additionally, CNN models demonstrate restricted predictive accuracy due to their dependence on convolutions if there are anomalies. Hu et al. [31] introduced the state deviation index for the diagnosis of FDI attacks and sudden load changes, demonstrating its effectiveness in IEEE 14-bus and 30-bus systems. However, the algorithm exhibited limited adaptability when faced with different errors. This suggests the need for an approach in which indicators are carefully designed to reflect unique properties of specific anomalies. The absence of such tailoring risks compromising the accuracy of evaluation metrics, presenting significant FP and false negative (FN) results. A multi-variable long short-term memory autoencoder (LSTM-AE) had been formulated for an SD process by Sarker et al. [32]. Their model successfully diagnosed SPs associated with errors, validated through a 123-bus unbalanced distribution network. However, according to sophisticated mathematical modeling of the PF analysis considering the SD process, developing the model with different abnormal scenarios in ML algorithms is challenging. Since they jeopardize the accuracy of the proposed algorithm, which can be time-consuming, they need more effort. Furthermore, a graphical user interface (GUI) of an EMS could show abnormal information due to system errors or bugs. These errors are not easy for SCADA control room engineers to diagnose due to the huge volume of information. These issues cannot be managed by the SE process and need adaptive solutions. A novel FDI diagnostic method using LSTM-AE and CNN-AE with an unsupervised learning approach was introduced in [33], avoiding the need for anomalous data during training. It also proposed an LSTM variational AE-based reconstruction method to maintain stability by closely replicating the original data from anomalous data. Despite this, the reliance on unsupervised learning presents challenges in certain scenarios. Hence, zero-day attacks need retraining of ML algorithms, which is time-intensive. Also, collecting all unknown errors can take much effort [34, 35]. Mukherjee [19] introduced an approach to identify FDI attacks using a multi-label classification framework. This method leveraged conventional bad-data detectors to enhance measurement accuracy and diagnose unstructured attacks. Accordingly, this model-free strategy required no prior knowledge of grid or attack vectors, making it a highly effective solution for the FDI diagnosis. Nevertheless, a consideration of an FDI attack cannot merely show the comprehensiveness of this algorithm. Also, there are some malfunctions that cannot be properly diagnosed by ML models. A retraining on new attacks and the diagnosis of other malfunctions (e.g., a CB can be opened during a fault in the normal operation; however, this could be because of inaccurate communication that sends the status of the CB to the SCADA room) are challenging for ML techniques [36, 37, 38]. Although recent studies have achieved notable advancements in utilizing LLMs for smart grids, there are still numerous areas that have not been fully addressed. Current methodologies either emphasize broad industrial frameworks that neglect grid-specific enhancements, concentrate solely on AD without predictive functions, or manage privacy with federated learning yet fail to achieve the necessary multimodal integration for robust smart grid security.
I-D Contributions
@@
  • Multi-Point Attack Detection Framework along with GenAI-based AD Processes: A GenAI-driven ADS model has been innovatively designed to transform the detection and resolution of vulnerabilities in the EMS data processing pipeline, targeting multi-point attacks and errors. In contrast to conventional approaches that focus on isolated detection points, this framework recognizes that sophisticated cyber-attacks exploit multiple stages of the SCADA data flow. The model specifically targets three critical vulnerability points including stealth attacks which can evade the BDDs, EMS database manipulation incidents (particularly FDI attacks following the SE process), and the HMI display corruption by manipulating the RTDB.
@@
  • Generative Intelligence-Enhanced Multimodal Analysis Framework: Addressing the integration challenges and anomaly identification requirements, an innovative SoM-GI framework is proposed. This contribution bridges the gap between traditional numerical SE techniques and optimal PF results and emerging visual AD capabilities. The framework exploits GenAI’s capabilities, augmented by visual markers and indicators alongside rules, to facilitate concurrent analysis of image and text data in the EMS environment. The SoM-GI methodology overcomes inherent spatial reasoning limitations in current GenAI systems by implementing systematic visual indicators, including CB status markers, directional transmission line indicators, and CP identifiers. These markers guide the AI’s interpretation of segmented HMI displays, enabling accurate detection of visual anomalies inconsistencies and falsifications that often bypass numerical detection methods. Validation results confirm the framework’s ability to identify sophisticated attacks that manipulate display segments while maintaining consistency in core data structures. This novel approach establishes a new paradigm for comprehensive security monitoring in power systems, integrating visual pattern recognition with linguistic rule processing. By standardizing the fusion of multimodal analysis techniques, the framework ensures that visual-based ADSs can be incorporated into existing EMS architectures without disrupting critical operational processes, thereby providing a practical solution.
I-E GenAI vs. Traditional ML/DL Approaches for Anomaly Detection in EMSs
The literature on the AD process in EMS domains has extensively explored classical ML and DL techniques. For example, unsupervised autoencoder frameworks reconstruct normal behavior and flag deviations via reconstruction error which in one recent industrial-substation case, the authors leveraged an LSTM-autoencoder and demonstrated that anomalies in remote-terminal sensing can be identified via elevated reconstruction losses [40]. In another study, graph neural network (GNN)-based methods contextualize measurement correlation across network nodes and show improved localization of cyber-injections compared to flat autoencoders [41]. Moreover, as reviewed in [42], many smart grid AD methods still rely on large volumes of labeled anomalies, assume stationary behavior, and are less capable of rapidly adapting to new attack modes.
In contrast, the proposed GenAI framework introduces several key differentiators. First, whereas conventional DL approaches often learn fixed mappings from feature space to latent representations (e.g., encoder–decoder), this generative model utilizes LLM-style embeddings and interactions to capture dynamic anomaly signatures and subtle deviations from normal. Second, by integrating a human-in-the-loop (HITL) feedback pathway, the system can incrementally refine its prompt embeddings in response to newly observed patterns, reducing the need for extensive labeled anomaly datasets. Third, this framework addresses zero-day or novel attack vectors using the generative synthesis of anomalous scenarios and continuous adaptation of the model’s understanding of normal versus abnormal behavior. Together, these capabilities position the GenAI approach as more flexible and future-proof for EMS/SCADA AD processes than traditional ML/DL baselines. Table II provides a systematic comparison of these approaches across key dimensions relevant to the EMS AD process, illustrating why GenAI represents a fundamental transformation rather than merely an incremental improvement over existing methodologies.
TABLE II: Comparison of GenAI vs. Traditional ML/DL Approaches for EMS Anomaly Detection [42].
Characteristic Traditional ML/DL Proposed GenAI Framework
Training Data Requirements Requires extensive labeled datasets covering all attack scenarios and operational conditions Leverages pre-trained knowledge; minimal domain-specific training data needed.
Feature Engineering Manual feature selection and engineering required; domain expertise intensive Automatic feature extraction from multimodal inputs; minimal manual engineering.
Adaptability Limited; requires complete retraining for new configurations or attack patterns High; few-shot and zero-shot learning enable rapid adaptation without retraining
Modality Processing Typically single-modality (numerical OR visual, not both) Native multimodal processing (numerical, visual, textual simultaneously)
Semantic Understanding Pattern recognition without semantic grounding in physical principles Deep semantic understanding of power system physics and operational constraints
Explainability Limited; black-box models with difficult-to-interpret outputs High; natural language explanations with specific rule/constraint violations cited
Zero-Day Attack Detection Poor; can only detect patterns seen during training Strong; can reason about novel attack patterns using physical principles
Domain Knowledge Integration Difficult; requires encoding expert knowledge into model architectures Natural; incorporates domain knowledge through natural language descriptions
Deployment Time Months (data collection, labeling, training, validation) Days to weeks (prompt engineering, rule specification, testing)
Visual Anomaly Detection Limited to computer vision models without operational context Integrated visual-semantic analysis with SoM-GI methodology
It is important to emphasize that GenAI does not necessarily replace traditional ML/DL approaches but rather complements them in a layered defense strategy. Conventional ML methods remain valuable for high-frequency, low-latency detection tasks where millisecond response times are critical. However, for complex reasoning tasks that require integration of multiple information sources, validation against physical constraints, interpretation of visual displays, and explanation of findings to human operators, GenAI offers capabilities that traditional approaches cannot match. This comparative analysis establishes the distinctive value proposition of the proposed GenAI-based framework and clarifies why it represents a significant advancement in EMS AD capabilities beyond what traditional ML/DL approaches can provide.
I-F Paper Structure